# ExplainReduce: Summarising local explanations via proxies

### v1

@@ 5.2 Fidelity of proxy sets
The interpretability of a set of local models as a global explanation is directly related to the size of that set. As explained in a previous section, local explanation models often need to balance interpretability and granularity: a large set of local models may more easily cover a larger portion of the closed-box function, but the increased number of models makes the result less understandable. In Fig. 5, we show the test fidelity (as defined in Eq. (8)) as a function of the size of the reduced explanation set for a selection of datasets and local explanation algorithms. The fidelity is calculated with respect to the closed-box predicted labels, with a fidelity of 0 representing perfect adherence to the closed-box model. As the figure shows, a set as small as $k=5$ proxies can reach or even surpass the fidelity of a set of $500$ local explanations for all three selected datasets***For full coverage on all of the seven datasets, we refer to Appendix B.. We find that optimisation-based approaches perform particularly well, especially the two greedy loss-minimising algorithms. The clustering-based reduction methods seem to work almost equally well, as does a random selection of local models, although optimisation-based approaches consistently outperform the latter. This implies that for proxy set generation methods that require choosing $k$ a priori, performance is not very sensitive to the choice of $k$. Therefore, it is enough for a user of ExplainReduce to specify a reasonable value of $k$ for a faithful and interpretable global surrogate.
@@
Figure 5: Test set fidelity of explanations as a function of proxy set size $k$. The rows show different datasets, and the columns show different XAI generation methods. The different line colours and styles denote the reduction strategy, and the black horizontal line shows the performance of the full explanation set. The loss-minimising reduction methods consistently reach a fidelity comparable or even better than the full explanation set.
@@
If a few proxies are enough to find a faithful representation, how many local explanations do we need to generate to find a good proxy set? The procedure may be infeasible if a very large set of initial local explanations is required. For example, slisemap scales as $\mathcal{O}(n^{2}m)$, where $n=|\mathcal{D}|$ and $m$ is the number of features [5], which makes generating thousands of explanations expensive.
@@
To study the generalisation performance of the proxy sets, we artificially limit the set of all local models $\bm{G}$ via random sampling before reduction. That is to say, we generate local approximations for each item in a randomly sampled subset of points in the training set ${\cal D}$. We then calculate the test fidelity of the proxy set trained on the limited universe as above and compare that to the performance of the full explanation set. As Fig. 6 shows, the proxy sets generalize well even starting from a limited set of local modelsFull version shown in Appendix C.. The coloured lines represent the various reduction algorithms, while the thick dashed black line denotes the performance of the entire set of local explanations. The loss-minimising greedy algorithms again achieve the best performance across the studied datasets and explanation methods studied, often matching and sometimes outperforming the full explanation set. On the other hand, the clustering-based approaches fall behind here, as in the previous section. These results imply that the procedure of ExplainReduce does not require training local models for each item; instead, a randomly sampled subset of models suffices to find a good proxy set. Based on these results, we set the default number of subsamples to $n=500$ in this paper unless otherwise stated.
@@
Figure 6: Test set fidelity of explanations as a function of local model subsample size $n=|\mathcal{D}|$. Different datasets are depicted as rows with the XAI methods used to produce the initial local explanations shown as rows. The coloured lines depict the test fidelity of the proxy sets produced with different reduction algorithms, with the thick black line denoting the performance of the entire local explanation set.
@@
We also experimented with the impact of the hyperparameters $\varepsilon$, the loss threshold, and $c$, the minimum coverage constraint, and found the reduction algorithms to have similar fidelity performance across a wide range of these hyperparameter values. We refer to Appendix F for detailed results.
5.3 Coverage and stability of the proxy sets
@@
It should be no surprise that methods directly optimising for loss also show the best fidelity results. However, an ideal global explanation method should be able to accurately explain most, if not all, of the items of interest. Additionally, explanations are expected to be locally consistent, meaning that similar data items should generally have similar explanations. In Fig. 7, we show the evolution of training coverage and instability of reduction methods as a function of the proxy set size $k$ with respect to the closed-box predicted labelsFull versions shown in Appendices D and E for coverage and stability, respectively.. The results follow a similar pattern to the previous sections: a small number of proxies is enough to achieve high coverage and low instability, with optimisation-based approaches outperforming clustering-based alternatives. Unsurprisingly, the coverage-optimising reduction methods achieve the best coverage with the smallest number of proxies. However, the loss-minimising algorithms perform surprisingly well even compared to the coverage-optimising variants. Notably, the balanced approach – the const min loss reduction algorithm – almost matches the reduction algorithms directly optimising coverage.
Regarding instability, the loss-minimising reduction methods have an edge over the clustering- and coverage-based approaches, but the distinction between methods is not pronounced.
Figure 7: Coverage and instability of various reduction methods as a function of $k$ on the Gas Turbine dataset. The optimisation-based methods outperform the clustering-based approaches within both metrics.
@@
Combining the results from the fidelity, coverage, and instability experiments, the greedy loss-minimising approach with a soft coverage constraint appears to offer the best balance in performance across all three metrics. The proxy set produced by this reduction method consistently achieves similar, if not better, performance than the full explanation set, even with a few proxies. Based on these experiments, the various clustering-based approaches do not offer performance gains based on any of the metrics used. While these approaches may produce more interpretable proxy sets by inherently accommodating the structures found in the data, they struggle to generate sufficiently covering proxy sets and lag behind optimisation-based approaches in both fidelity and instability.
Coverage $\uparrow$
Dataset G. Max $c$ A. ratio G. Min Loss (min $c$) A. ratio Max $c$
Air Quality $0.82\pm 0.18$ $0.97\pm 0.04$ $0.9\pm 0.05$ $0.93\pm 0.03$ $0.84\pm 0.18$
Gas Turbine $0.86\pm 0.17$ $0.99\pm 0.01$ $0.93\pm 0.06$ $0.97\pm 0.03$ $0.87\pm 0.18$
Jets $0.86\pm 0.17$ $0.99\pm 0.02$ $0.95\pm 0.06$ $0.98\pm 0.03$ $0.87\pm 0.18$
QM9 $0.78\pm 0.17$ $0.98\pm 0.03$ $0.89\pm 0.04$ $0.96\pm 0.02$ $0.8\pm 0.18$
Train fidelity $\downarrow$
Dataset G. Min Loss A. ratio G. Min Loss (min $c$) A. ratio Min Loss
Air Quality $0.36\pm 0.35$ $1.25\pm 0.21$ $0.16\pm 0.12$ $1.53\pm 0.53$ $0.32\pm 0.35$
Gas Turbine $0.21\pm 0.23$ $2.08\pm 0.9$ $0.07\pm 0.07$ $1.53\pm 0.43$ $0.17\pm 0.23$
Jets $0.02\pm 0.02$ $1.4\pm 0.4$ $0.01\pm 0.0$ $1.41\pm 0.27$ $0.01\pm 0.02$
QM9 $0.53\pm 0.46$ $1.25\pm 0.21$ $0.24\pm 0.15$ $1.37\pm 0.24$ $0.48\pm 0.47$
Test fidelity $\downarrow$
Dataset G. Min Loss A. ratio G. Min Loss (min $c$) A. ratio Min Loss
Air Quality $0.4\pm 0.29$ $1.11\pm 0.22$ $0.22\pm 0.07$ $1.15\pm 0.17$ $0.37\pm 0.28$
Gas Turbine $0.31\pm 0.17$ $1.2\pm 0.22$ $0.21\pm 0.06$ $1.1\pm 0.11$ $0.28\pm 0.18$
Jets $0.02\pm 0.01$ $1.09\pm 0.18$ $0.01\pm 0.0$ $1.04\pm 0.1$ $0.02\pm 0.01$
QM9 $1.19\pm 0.13$ $1.01\pm 0.08$ $1.09\pm 0.06$ $0.96\pm 0.09$ $1.18\pm 0.11$
Table 2: Comparison of the training fidelity and coverage, alongside the observed approximation ratios (A. ratio), between the analytical and approximate optimisation algorithms when applied to $n=100$ explanations produced with lime and with $k=5$ proxies. G. Max $c$ corresponds to the greedy max coverage algorithm, while G. Min Loss and G. Min Loss (min $c$) correspond to the greedy min loss and const min loss algorithms, respectively.
5.4 Performance of greedy algorithms
In section 4.3, we discussed the performance of greedy approximation algorithms used to solve Problems 13. As we have seen in previous sections, the performance of the greedy and analytical coverage-optimising algorithms is almost equal as both a function of the proxy set size and the initial local explanation set size. As mentioned in the section above, solving for the exact loss-minimising proxy set is a computationally challenging problem which requires $\binom{m}{k}$ comparisons in the worst case. Hence, the analysis in this section was limited to $n=m=100,k=5$ for computational tractability. Table 2 shows how the greedy approximation algorithms perform comparably to the analytically optimal variants. For the max coverage algorithm, experimental approximation ratios are substantially better than the worst-case bound described in Sect. 4.3. Furthermore, as we saw in the previous sections, while on the train set the greedy heuristics do not reach the optimal solution fidelity-wise, on the test the differences between the exact solution and the greedy heuristics vanish. The const min loss algorithm especially shows poorer performance compared to the greedy min loss algorithm but generalises better on the test set while achieving good coverage performance as well. Overall, we can conclude that the greedy approximations offer a computationally feasible approach to generating proxy sets.
@@ 6 Discussion
Using local surrogate models as explanations is a common approach for model-agnostic XAI methods. In this paper, we have shown that the ExplainReduce procedure can find a post-hoc global explanation consisting of a small subset of simple models when provided a modest set of pre-trained local explanations. The procedure is agnostic to both the underlying closed-box model and the XAI method, assuming that the XAI method can function as a generative (predictive) model. These proxy sets achieve comparable fidelity, stability, and coverage to the full explanation set while remaining succinct enough to be easily interpretable by humans. The loss-minimising reduction method with a soft coverage constraint especially offers a balance of fidelity, coverage, and stability across different datasets. Furthermore, we have shown how we can use greedy algorithms to find these proxy sets with minimal computational overhead with nearly equal performance to analytically optimal counterparts. The methods are not sensitive to hyperparameters, allowing for easy adoption.
As with any work, there are shortcomings. The presented implementation for the procedure simply finds a set of proxies that is (approximately) optimal from a global loss or coverage perspective. In other words, there is no consideration of the spatial distribution of the data with respect to which proxy was associated with each item. To deepen the insight provided by the method, an interesting idea would be to combine the reduction process with manifold learning techniques, enabling visualisation alongside the proxy set, similar to slisemap and slipmap. For example, we might try to find an embedding where data items associated with a particular proxy would form continuous regions, significantly increasing interpretability and augmenting the method’s capabilities as an exploratory data analysis tool.
Another interesting way to leverage the set of applicable local models is uncertainty quantification. Instead of trying to find local models applicable to many data items, we could instead turn our attention to all applicable models for a given item. This set of applicable models could be considered as samples from the global distribution of explanations for that particular item. Thus, we could use such sets (alongside the produced explanation) as a way to find, e.g., confidence intervals for explanations produced by a wide variety of XAI methods.
@@ Acknowledgements
Appendix A Dataset introduction
@@
Synthetic [5] is a synthetic regression dataset with $N$ data points, $M$ features, $k$ clusters, and cluster spread $s$. Each cluster is associated with a linear regression model and a centroid. Coefficient vectors $\beta_{j}\in\mathbb{R}^{M+1}$ and centroids $c_{j}\in\mathbb{R}^{M}$ are sampled from normal distributions, with repeated samplings if either the coefficients or the centroids are too similar between the clusters. Each data point is assigned to a cluster $j_{i}$, with features $x_{i}$ sampled around $c_{j_{i}}$ and standardized. The target variable is $y_{i}=x_{i}^{\top}\beta_{j_{i}}+\epsilon_{i}$, where $\epsilon_{i}$ is Gaussian noise with standard deviation $\sigma_{e}$. In the experiment section, unless otherwise stated, we set the number of data points to $N=5000$, the number of features to $M=11$, the number of clusters to $k=5$, and the standard deviation of Gaussian noise to $\sigma_{e}=2.0$. As the closed-box model, we use a random forest regressor.
@@
Air Quality [28] contains 7355 hourly observations of 12 different air quality measurements. One of the measured qualities is chosen as the label, and the other values are used a covariates. With this dataset, we use a random forest regressor as the closed-box model.
@@
Life Expectancy [29] comprising 2,938 instances of 22 distinct health-related measurements, this dataset spans the years 2000 to 2015 across 193 countries. The expected lifespan from birth in years is used as the dependent variable, while other features act as covariates. The closed-box model we use with this dataset is a neural network.
@@
Vehicle [30] contains 2059 instances of 12 different car-related features, with the target being the resale value of the instance. We use a support vector regressor as the closed-box model with the Vehicles dataset.
@@
Gas Turbine [31] is a regression dataset with 36,733 instances of 9 sensor measurements on a gas turbine to study gas emissions. With this dataset, we used a Adaboost regression as the closed-box model.
@@
QM9 [32] is a regression dataset comprising 133,766 small organic molecules. Features are created with the Mordred molecular description calculator [36]. The QM9 closed-box model is a neural network.
@@
HIGGS [34] is a two-class classification dataset consisting of signal processes that produce Higgs bosons or are background. The dataset contains nearly $100000$ instances with 28 features. We used a gradient boosting classifier with this dataset as the closed-box model.
@@
Jets [33] contains simulated LHC proton-proton collisions. The collisions produce quarks and gluons that decay into cascades of stable particles called jets. The classification task is to distinguish between jets generated by quarks and gluons. The dataset has $266421$ instances with 7 features. The closed-box model we used with this dataset is a random forest classifier.
@@
Appendix B Fidelity of the proxy sets as a function of proxy number
Figure 8 shows the complete fidelity plots as a function of $k$. Each row corresponds to a different dataset, and each column represents a local model generation method. A general trend emerges, indicating improved fidelity with increasing $k$.
Figure 8: Fidelity of the reduction algorithms as a function of the proxy set size $k$. The black dashed line indicates the performance of the full set of explanations.
Appendix C Fidelity of the proxy sets as a function of subsample size

### v2

@@ 5.2.1 Fidelity of proxy sets
The interpretability of a set of local models as a global explanation can be directly linked to the size of that set. As explained in a previous section, local explanation models often need to balance interpretability and granularity: a large set of local models may more easily cover a larger portion of the closed-box function, but the increased number of models makes it less understandable. In Fig. 5, [^2] For full coverage on all of the twelve datasets, we refer to Appendix C. we show the test fidelity (as defined in Eq. (9)) as a function of the size of the reduced explanation set for a selection of datasets and local explanation algorithms, following the naming convention presented in Section 4.1. The fidelity is calculated with respect to the closed-box predicted labels, with a fidelity of zero representing perfect adherence to the closed-box model. As the figure shows, a set as small as $k=5$ proxies can reach or even surpass the fidelity of a set of $500$ local explanations for the selected datasets (horizontal black dashed line). As expected, the two loss-minimising algorithms (green line for Min Loss and red for Balanced) have lower fidelity than the greedy coverage-maximising algorithm (blue line). All three algorithms outperform the baseline of randomly picking explanations (yellow dotted line). We also note that while the fidelity initially drops rapidly as $k$ increases, the metric quickly plateaus beyond $k>5$. Therefore, it is enough for a user of ExplainReduce to specify a reasonable value of $k$ for a faithful and interpretable global surrogate.
@@
Figure 5: As few as $k=5$ proxies are able to reach nearly or better test set fidelity than the full set of local explanations. Lower values are better. The rows show different datasets, and the columns show different XAI generation methods. The different line colours and styles denote the reduction strategy, and the black horizontal line shows the performance of the full explanation set. The loss-minimising reduction methods, in particular, consistently reach a fidelity comparable to or even better than the full explanation set. Smaller fidelity is better.
@@
If a few proxies are enough to find a faithful representation, how many local explanations do we need to generate to find a good proxy set among them? After all, the procedure may be infeasible if a very large set of initial local explanations is required. For example, sampling-based XAI methods (lime, shap, and smoothgrad) scale as $\mathcal{O}(nm)$, where $n=|\mathcal{D}|$ and $m$ is the number of samples per local model, which makes generating thousands of explanations expensive.
@@
To study the effect of local model number to proxy set performance, we artificially limit the set of all local models $\bm{G}$ via random sampling between generation and reduction. We then calculate the test fidelity of the proxy set trained on the limited universe, as described above, and compare that to the performance of the full explanation set. As Fig. 6 shows, the proxy sets generalise well even starting from a limited set of local models [^3] Full version shown in Appendix D.. The coloured lines represent the various reduction algorithms, using the same colouring convention as in Fig. 5, while the thick dashed black line denotes the performance of the initial set of local explanations. The loss-minimising greedy algorithms again achieve the best performance across the studied datasets and explanation methods, often matching and sometimes outperforming the full explanation set. These results imply that ExplainReduce does not require training local models for each item; instead, a randomly sampled subset of models suffices to find a good proxy set. Based on these results, we set the default number of subsamples to $n=500$ in the remainder of this paper.
@@
Figure 6: Test set fidelity quickly plateaus as the size of the initial local explanation set $n=|\mathcal{\bm{G}}|$ increases. Different datasets are depicted as rows, with the XAI methods used to produce the initial local explanations shown as rows. The coloured lines depict the test fidelity of the proxy sets produced with different reduction algorithms, with the thick black line denoting the performance of the entire local explanation set. Smaller fidelity is better.
@@
We also experimented with the impact of the hyperparameters $\varepsilon$, the loss threshold, and $\lambda$, the trade-off parameter. We found that the reduction algorithms maintain similar fidelity performance across a wide range of these hyperparameter values. We refer to Appendix E for detailed results.
5.2.2 Coverage and stability
@@
It should be no surprise that methods directly optimising for loss also show the best fidelity results. However, an ideal global explanation method should be able to accurately explain most, if not all, of the items of interest. Additionally, explanations are expected to be locally consistent, meaning that similar data items should generally have similar explanations. In Fig. 7, we show the evolution of training coverage and instability of reduction methods as a function of the proxy set size $k$ with respect to the closed-box predicted labels [^4] Full versions shown in Appendix C.. The results follow a similar pattern to the previous sections: a small number of proxies is enough to achieve high coverage and low instability, with optimisation-based approaches outperforming clustering-based alternatives. Unsurprisingly, the coverage-optimising reduction method achieves the best coverage with the smallest number of proxies. However, the loss-minimising algorithms perform surprisingly well, even compared to the coverage-optimising variant. Notably, the balanced approach—the Balanced reduction algorithm—almost matches the reduction algorithm directly optimising coverage.
The instability results show a similar trend to the fidelity results: the loss-minimising approaches have a slight edge over the coverage-maximising approach.
Figure 7: Small proxy sets also reach similar coverage and instability values as the full local explanation set. These results are from the Gas Turbine regression dataset. Naturally, the coverage-maximising algorithm shows best performance for coverage, but the balanced Balanced algorithm has nearly equal results. Larger coverage and smaller instability are better.
@@
Combining the results from the fidelity, coverage, and instability experiments, the greedy Balanced reduction algorithm appears to offer the best balance in performance across all three metrics. The proxy set produced by this reduction method consistently achieves performance similar to, if not better than, the full explanation set, even with a few proxies. Furthermore, these results and those presented in Appendix G demonstrate more favourable practical approximation ratios than the worst-case performance discussed in Section 4.1 for all three algorithms.
5.3 Other global explanation and explanation aggregation methods
To verify the performance of the best ExplainReduce algorithm, the Balanced reduction, we compared the method to submodular pick of lime, GLocalX as well as the IP-based approach by [li2022optimal] (here termed max spherical coverage due to their definition of coverage through spherical neighbourhoods). Note, however, that GLocalX and the IP-based approach are limited to classification problems, and the submodular pick algorithm is limited to lime only, unlike ExplainReduce, which applies to the full gamut of tasks and XAI methods. In Figure 8555Results for other datasets available in Appendix F., we show the train and test fidelities and the train coverage for the model aggregation methods on the Adult dataset. ExplainReduce shows better or matching fidelity on both the train and test sets, showing that the method is able to find the most faithful subset of explanations both from the initial set of local models and also when generalising to novel items. Similarly, our method achieves the highest or matching train set coverage, meaning that the subset found by our method has the highest likelihood on containing a reasonable local model for a given item. It should also be noted that while ExplainReduce and submodular pick are computationally very efficient and find the proxy set in a matter of seconds, the IP-based max spherical coverage and GLocalX algorithms take hundreds or thousands of seconds for the same task on the same hardware. For some datasets, such as the Adult dataset depicted here, a single global linear model (horizontal red dashed line) has lower test set fidelity than even the full XAI local model sets with about $60\%$ coverage. The local model sets and reductions outperform the single global model in train fidelity and coverage, meaning that more items in the test set can be associated with a local explanation within the set error tolerance. As the rules produced by lore are weaker predictors due to their simplicity, all the reduction methods are able to match the test set performance of the full rule set with very few rules. Hence, we cannot say much about the relative performance of the methods beyond noting that ExplainReduce generally performs comparatively to the rule-specific GLocalX.
Figure 8: ExplainReduce, using the Balanced reduction algorithm (green dashed line), is able to match or exceed test set fidelity of other model aggregation methods. The black dashed line shows the performance of all local explanations. The red horizontal dashed line shows the fidelity of a single sparse global linear model, which, however, only covers about 60% of the datapoints (bottom row). Smaller fidelity and higher coverage is better.
@@ 6 Discussion
Using local surrogate models as explanations is a common approach for model-agnostic XAI methods. In this paper, we have shown many local explanations are redundant, and that the ExplainReduce procedure can find a post-hoc global explanation consisting of a small subset of simple models when provided a modest set of pre-trained local explanations. The procedure is agnostic to both the underlying closed-box model and the XAI method, assuming that the XAI method can function as a generative (predictive) model. These proxy sets achieve comparable fidelity, stability, and coverage to the full explanation set while remaining succinct enough to be easily interpretable by humans. The Balanced algorithm especially offers good performance in fidelity, coverage, and stability across different datasets. Furthermore, we have shown how our approach shows favourable performance over other global explanation methods. The procedure is not sensitive to hyperparameters, allowing for easy adoption.
The main limitation of our approach is that the presented implementation for the procedure simply finds a set of proxies that is (approximately) optimal from a global loss or coverage perspective. In other words, there is no consideration of the spatial distribution of the data with respect to which proxy was associated with each item. To deepen the insight provided by the method, an interesting idea would be to combine the reduction process with manifold learning techniques, enabling visualisation alongside the proxy set, similar to slisemap and slipmap. For example, we might try to find an embedding where data items associated with a particular proxy would form continuous regions, significantly increasing interpretability and augmenting the method’s capabilities as an exploratory data analysis tool.
Another interesting way to leverage the set of applicable local models is uncertainty quantification. Instead of trying to find local models applicable to many data items, we could focus on all applicable models for a given item. This set of applicable models could be considered as samples from the global distribution of explanations for that particular item. Thus, we could use such sets (alongside the produced explanation) as a way to find, e.g., confidence intervals for explanations produced by a wide variety of XAI methods.
@@ Acknowledgements
Appendix A Details on XAI methods used
In this section, we give an overview of the formulation of XAI methods analysed in this article.
Theoretically, the simplest way to produce a local surrogate model would be to calculate the gradient of the closed-box function. This XAI method is often called vanillagrad in the literature. However, in practice, the gradients of machine learning models can be very noisy, as demonstrated by the effectiveness of adversarial attacks that exploit small perturbations in neural networks [szegedy2014intriguing]. On the other hand, many commonly used machine learning models, such as random forests, do not have well-defined gradients. Hence, more involved approaches are warranted.
smoothgrad [smilkov2017smoothgrad] attempts to solve the gradient noise problem in vanillagrad by averaging over gradients sampled from the vicinity of the point under explanation. Although the original paper only applies smoothgrad to classification, the method can easily be extended to regression.
Moving away from directly analysing the gradient, lime [ribeiro2016] and shap [lundberg2017unified] are perhaps the most widely known examples of practical local explanation generation methods. In this paper, we focus on the kernel-shap variant, which combines linear lime with shap. Both lime and kernel-shap produce explanations in the form of additive linear models ($\tilde{y}=\phi^{T}\bm{x}$). Given an item $\bm{x}_{i}$, the methods sample novel items $\bm{x}^{\prime}_{i}\in\mathcal{X}^{\prime}$ in the neighbourhood of the first item, use the closed-box model to predict the labels for the novel items and find the linear model that best fits them [lundberg2017unified],
where $\pi_{x}$ represents a distance measure and $\Omega(\phi)$ is a regularisation term. The difference between the methods lies in the choice of distance measure, which defines the notion of neighbourhood for $\bm{x}_{i}$; lime most often uses either $L^{2}$ or cosine distance, whereas kernel-shap utilises results in game theory [slack2021reliable].
smoothgrad, lime and shap are based on sampling novel items, which, while simple to implement, introduce a unique set of challenges. First, formulating a reliable data generation process or sampling scheme for all possible datasets is difficult, if not impossible [guidotti2018survey, laugel2018defining]. For example, images generated by adding random noise to the pixel values rarely resemble natural images. Second, randomly generating new items might produce items that cannot occur naturally due to, for example, violating the laws of physics. slisemap [bjorklund2023slisemap] and its variant, slipmap [bjorklund2024SLIPMAP], produce both a low-dimensional embedding for visualisation and a local model for all training items without sampling any new points. slisemap finds both the embedding and local models by optimising a loss function consisting of two parts: an embedding term, where items with similar local explanations attract each other while repelling dissimilar ones and a local loss term for the explanation of each item:
$\min_{g_{i}}\mathcal{L}_{i}=\sum\nolimits_{i=1}^{n}\frac{\exp(-D(\bm{z}_{i},\bm{z}_{j}))}{\sum_{k=1}^{n}\exp(-D(\bm{z}_{k},\bm{z}_{j}))}\ell(g_{i}(\bm{x}_{j}),\bm{\hat{y}}_{j})+\Omega(g_{i})$ (13)
As an example of rule-based explanation method (to be paired with the GLocalX aggregation method) we consider Local Rule-based Explanations (lore) [guidotti2019lore]. Rule-based local explanations take the form of simple decision rules, such as $r=\{age\leq 25;job=clerk;income\leq 900\}\rightarrow deny$ for a loan application classification task. In lore, these rules are generated via sampling the neighbours of a given item using a genetic algorithm to explore the decision boundary around the item. Then, the closed-box model is used as an oracle for the generated neighbourhood items data, and a decision tree is trained with the predicted labels, from which the resulting rule is derived. The procedure is hence similar to generating lime and kernel-shap explanations. The rules learned from the training data form a rule set. To use a rule set to predict novel items, we followed the procedure used by the authors of the article proposing GLocalX [setzu2021glocalx] and outlined by [yin2003cpar]. We first find all rules applicable to a given novel item, and then use the rule with the highest accuracy on the test set among these to produce the prediction.
Appendix B Dataset introduction
@@
Synthetic [bjorklund2023slisemap] is a synthetic regression dataset with $N$ data points, $M$ features, $k$ clusters, and cluster spread $s$. Each cluster is associated with a linear regression model and a centroid. Coefficient vectors $\beta_{j}\in\mathbb{R}^{M+1}$ and centroids $c_{j}\in\mathbb{R}^{M}$ are sampled from normal distributions, with repeated samplings if either the coefficients or the centroids are too similar between the clusters. Each data point is assigned to a cluster $j_{i}$, with features $x_{i}$ sampled around $c_{j_{i}}$ and standardised. The target variable is $y_{i}=x_{i}^{\top}\beta_{j_{i}}+\epsilon_{i}$, where $\epsilon_{i}$ is Gaussian noise with standard deviation $\sigma_{e}$. In the experiment section, unless otherwise stated, we set the number of data points to $N=5000$, the number of features to $M=11$, the number of clusters to $k=5$, and the standard deviation of Gaussian noise to $\sigma_{e}=2.0$. In the closed-box model, we use a random forest regressor.
@@
Air Quality [oikarinenDetectingVirtualConcept2021] contains 7355 hourly observations of 12 different air quality measurements. One of the measured qualities is chosen as the label, and the other values are used as covariates. With this dataset, we use a random forest regressor as the closed-box model.
@@
Life Expectancy [rajarshi2017life] comprises 2,938 instances of 22 distinct health-related measurements, spanning the years 2000 to 2015 across 193 countries. The expected lifespan from birth in years is used as the dependent variable, while other features act as covariates. The closed-box model we use with this dataset is a neural network.
@@
Vehicle [birla2022vehicles] contains 2059 instances of 12 different car-related features, with the target being the resale value of the instance. We use a support vector regressor as the closed-box model with the Vehicles dataset.
@@
Gas Turbine [2019gas] is a regression dataset with 36,733 instances of 9 sensor measurements on a gas turbine to study gas emissions. With this dataset, we used an Adaboost regression as the closed-box model.
@@
QM9 [ramakrishnan2014Quantum] is a regression dataset comprising 133,766 small organic molecules. Features are created with the Mordred molecular description calculator.The QM9 closed-box model is a neural network.
Adult [becker1996adult] is a classic classification dataset with 48,842 observations of individuals and 14 features based on census data. The goal is to predict whether or not an individual’s annual income exceeds \$50,000. The closed-box model we used was a neural network classifier.
Churn [blake1998churn] is a classification dataset with 5,000 instances and 20 features. The goal is to predict whether a customer of a telecommunications company left the services of that company (“churned”) using features that represent information about the customer and their telephone use. With this dataset, we used a random forest closed-box model.
@@
HIGGS [whiteson2014higgs] is a two-class classification dataset consisting of signal processes that produce Higgs bosons or are background. The dataset contains nearly $100000$ instances with 28 features. We used a gradient boosting classifier with this dataset as the closed-box model.
@@
Jets [CMS:opendata] contains simulated LHC proton-proton collisions. The collisions produce quarks and gluons that decay into cascades of stable particles called jets. The classification task is to distinguish between jets generated by quarks and gluons. The dataset has $266421$ instances with 7 features. The closed-box model we used with this dataset is a random forest classifier.
Spam [hopkins1999spam] or Spambase is a classification dataset with 4,601 instances and 57 features. The dataset consists of various email metadata, and the labels correspond to whether or not the email is spam. With this dataset, we used a gradient boosting classifier as the closed-box model.
Telescope [bock2004telescope] is a classification dataset with 13,376 instances with 10 features. The data are Monte Carlo simulated instances of Cherenkov radiation produced by high-energy gamma rays hitting the atmosphere, as they would be detected by a ground-based atmospheric Cherenkov gamma telescope. The goal is to classify high-energy gamma events from background observations. The closed-box model we used with this dataset was a neural network classifier.
@@
Appendix C Fidelity, coverage and stability of the proxy sets as a function of proxy number
Figures 9-20 show the complete fidelity, coverage and stability plots as a function of $k$. Each figure corresponds to a different dataset, and each column represents a local model generation method. ExplainReduce achieves strong generalisation across all datasets. The loss-minimisation algorithms outperform or match coverage-optimising approaches, while remaining computationally efficient. Although slisemap often provides the most stable local explanations, the results indicate that the reduction algorithm—rather than the explainer alone—determines the quality of the global surrogate. The combination of slisemap with the Greedy Minimum loss reduction method consistently yields low test-set loss and high coverage. In most datasets, as few as five proxy models ($k\approx 5$) suffice to approximate or even surpass the fidelity of the full ensemble of local explanations. These results confirm that ExplainReduce can drastically compress hundreds of local models into a handful of interpretable surrogates while retaining predictive consistency.
Figure 9: Adult dataset reduction function comparison as a function of $k$ on fidelity, coverage and stability.
Figure 10: Air Quality dataset reduction function comparison as a function of $k$ on fidelity, coverage and stability.
Figure 11: Churn dataset reduction function comparison as a function of $k$ on fidelity, coverage and stability.
Figure 12: Gas Turbine dataset reduction function comparison as a function of $k$ on fidelity, coverage and stability.
Figure 13: Higgs dataset reduction function comparison as a function of $k$ on fidelity, coverage and stability.
Figure 14: Jets dataset reduction function comparison as a function of $k$ on fidelity, coverage and stability.
Figure 15: Life Expectancy dataset reduction function comparison as a function of $k$ on fidelity, coverage and stability.
Figure 16: QM9 dataset reduction function comparison as a function of $k$ on fidelity, coverage and stability.
Figure 17: Spam dataset reduction function comparison as a function of $k$ on fidelity, coverage and stability.
Figure 18: Synthetic dataset reduction function comparison as a function of $k$ on fidelity, coverage and stability.
Figure 19: Telescope dataset reduction function comparison as a function of $k$ on fidelity, coverage and stability.
Figure 20: Vehicles dataset reduction function comparison as a function of $k$ on fidelity, coverage and stability.
Appendix D Fidelity of the proxy sets as a function of subsample size